OpenStudy (anonymous):

in a 1000 m race, A beats B by 50m or 10 sec find speed of A

OpenStudy (anonymous):

@Directrix @thomaster

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@BarbaraKara

OpenStudy (anonymous):

speed of b = y m/s speed of a = x m/s 1000m = at 1000m - 50m = b(t - 10) -> 1000m = b(t-10) + 50m so, at = b(t - 10) + 50m so a = [b(t - 10) + 50m] / t if u have the values of b and t, u get the answer straight away..

OpenStudy (anonymous):

there are no values for b and t

OpenStudy (anonymous):

umm. then i think that's all that i can help.. i'm limited..

OpenStudy (anonymous):

k thanks